# High-Dimensional Bayesian Model Comparison in Cosmology with GPU-accelerated Nested Sampling and Neural Emulators

### v1

@@ Acknowledgements
Appendix A Cosmic Shear Likelihood
We consider a tomographic survey with $N_{\text{bins}}=10$ redshift bins. The primary observable is the angular power spectrum of the cosmic shear signal $C^{\epsilon\epsilon}_{ij}(\ell)$ between all pairs of bins $(i,j)$. This signal is composed of the true cosmic shear signal ($\gamma$) and a significant contaminant from the intrinsic alignments ($I$) of galaxies:
@@ Appendix A Cosmic Shear Likelihood
Each of these components is computed using the extended Limber approximation (LoVerde and Afshordi, 2008), which projects the 3D matter power spectrum, $P_{\delta\delta}(k,z)$, into a 2D angular correlation:
@@
which is sensitive to the matter density $\Omega_{m}$, the Hubble parameter $H_{0}$ and the scale factor $a$. To achieve the precision required by modern surveys, we must also model several systematic effects, each of which introduces nuisance parameters that must be marginalized over.
  • Intrinsic Alignments ($A_{\text{IA},i}$): The intrinsic shapes of nearby galaxies can be physically correlated, mimicking a lensing signal. We model this using the Non-Linear Alignment (NLA) model (Hirata and Seljak, 2004) as parametrised in Piras and Spurio Mancini (2023), which defines the intrinsic alignment window function $W^{I}$. We introduce a single free amplitude parameter per bin, $A_{\text{IA},i}$, controlling the strength of the signal:
@@
    which is sensitive to the linear growth factor $D(\chi)$, the critical density of the universe $\rho_{\rm cr}$, and a normalization constant $C_{1}$ fixed to $5\times 10^{-14}\,h^{-2}M_{\odot}^{-1}\text{Mpc}^{3}$. This adds 10 nuisance parameters to our model.
@@
  • Photometric Redshift Errors ($D_{z_{i}}$): Uncertainties in photometric redshift estimation can shift galaxies between bins. We model this with a simple shift parameter $D_{z_{i}}$ for the mean of each redshift distribution, $n_{i}(z)\rightarrow n_{i}(z-D_{z_{i}})$ (Eifler et al., 2021), adding a final 10 nuisance parameters.
The likelihood for the data vector of all angular power spectra is assumed to be a multivariate Gaussian, with a simulated covariance matrix for an LSST-like survey ($n_{\text{gal}}=30\,\text{arcmin}^{-2}$, $\sigma_{\epsilon}=0.3$, $f_{\text{sky}}=0.35$).
The three likelihoods mentioned in table 2: Shear, Shear (I), and Shear (*) are differentiated by their exact implementation of the above. (*) is the original jax_cosmo implementation of the above, as used in (Piras et al., 2024). The main shear likelihood of this paper is the re-factored version of (*) that aimed to be a more JAX-native and optimised version of that code, while Shear (I) is the interpolated version of this optimised likelihood. (I) trades some accuracy by evaluating the matter power spectrum at a grid at 10 times lower resolution than the other likelihoods and then interpolates over $z$ and then the $k$-modes.
Appendix B Full Cosmic Shear Constraints

### v2

@@ Acknowledgements
Appendix A Cosmic Shear Likelihood Details
We consider a tomographic survey with $N_{\text{bins}}=10$ redshift bins. The primary observable is the angular power spectrum of the cosmic shear signal $C^{\epsilon\epsilon}_{ij}(\ell)$ between all pairs of bins $(i,j)$. This signal is composed of the true cosmic shear signal ($\gamma$) and a contaminant from the intrinsic alignments ($I$) of galaxies:
@@ Appendix A Cosmic Shear Likelihood Details
Each of these components is computed using the extended Limber approximation (LoVerde and Afshordi,, 2008), which projects the 3D matter power spectrum, $P_{\delta\delta}(k,z)$, into a 2D angular correlation:
@@
which is sensitive to the matter density $\omega_{\rm{m}}$, the Hubble parameter $H_{0}$ and the scale factor $a$. To achieve the precision required by modern surveys, we must also model several systematic effects, each of which introduces nuisance parameters that must be marginalised over.
  • Intrinsic Alignments ($A_{\text{IA},i}$): The intrinsic shapes of nearby galaxies can be physically correlated, mimicking a lensing signal. We model this using the Non-Linear Alignment (NLA) model (Hirata and Seljak,, 2004) as parametrised in Piras and Spurio Mancini, (2023), which defines the intrinsic alignment window function $W^{I}$. We introduce a single free amplitude parameter per bin, $A_{\text{IA},i}$, controlling the strength of the signal:
@@
    which is sensitive to the linear growth factor $D(\chi)$, the critical density of the universe $\rho_{\rm cr}$, and a normalisation constant $C_{1}$ fixed to $5\times 10^{-14}\,h^{-2}M_{\odot}^{-1}\text{Mpc}^{3}$. This adds 10 nuisance parameters to our model.
@@
  • Photometric Redshift Errors ($D_{z_{i}}$): Uncertainties in photometric redshift estimation can shift galaxies between bins. We model this with a simple shift parameter $D_{z_{i}}$ for the mean of each redshift distribution, $n_{i}(z)\rightarrow n_{i}(z-D_{z_{i}})$ (Eifler et al.,, 2021), adding a final 10 nuisance parameters.
The likelihood for the data vector of all angular power spectra is assumed to be a multivariate Gaussian, with a simulated covariance matrix for an LSST-like survey ($n_{\text{gal}}=30\,\text{arcmin}^{-2}$, $\sigma_{\epsilon}=0.3$, $f_{\text{sky}}=0.35$). The $C_{ij}^{AB}$ and $W$ integrals are all evaluated on a grid of 10251025 $\chi$ values that are evenly spaced in scale factor from $a=1$ down to the furthest object in the mock data, $a_{\text{min}}=(1+5.02)^{-1}=0.16$.
The three likelihoods mentioned in Table 2, namely Shear, Shear (I), and Shear (*), are differentiated by their exact implementation of the $C^{\epsilon\epsilon}_{ij}(\ell)$. (*) is the original jax_cosmo implementation of the likelihood as used in Piras et al., (2024). The main shear likelihood of this paper is the refactored version of (*) that aimed to be a more JAX-native and optimised version of that code, while Shear (I) is the interpolated version of this optimised likelihood.
Appendix B Accuracy of Interpolated Likelihoods